OpenStudy (anonymous):

Simplify completely quantity 4 x plus 16 over 6 I got 2 open parentheses x plus 4 close parentheses over 3

OpenStudy (ash2326):

\[4\times \frac{16}{6}\] Is this your question?

OpenStudy (ash2326):

or \[\frac{4x+16}{6}\] ??

OpenStudy (anonymous):

no its 4x + 16 divided by 6

OpenStudy (anonymous):

the second one

OpenStudy (ash2326):

Take 2 common from both the numerator and the denominator \[\frac{2(2x+8)}{2(3)}\] Cancel the 2s \[\frac{\cancel 2(2x+8)}{\cancel2(3)}=>\frac{2x+8}{3}\]

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

2(x+8) over 3 is wrong for future people